# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v1

@@ 1 Introduction
Here, based on MinAP [48, 49], we formulate the generalized finite-time optimal control (GFTOC) framework valid for any discrete-state fEQ system, by exploiting the Riemannian geometric structure in the control parameter space. We compute the optimal finite-time driving protocols that minimize the total driving EP, which exhibit the ‘kinks’. GFTOC unifies the optimal control of slowly driven and finite-time processes, as we prove an exact mapping between them. Importantly, since MinAP assigns a thermodynamic cost for any transition path realization, a thermodynamic cost for sustaining the ‘kinks’ is quantified (a boundary term in driving time), namely ‘thermodynamic shock’. Thereby, the ‘kinks’ reduce the thermodynamic cost of driving (a bulk term in driving time), since the ‘kinks’ reduce the distance to be travelled in the control parameter space. The thermodynamic interplay between the boundary and the bulk terms of driving EP gives rise to the formation of the ‘kinks’ as a dynamic consequence. The dual manifestation of the bulk-boundary term of EP driving has a physical interpretation analogous to the work-heat, but formulated here for the finite-time optimal process in ST. We show that the dynamic counterpart of this thermodynamic phenomena is dictated by the restoration of the ‘finite-time speed limit’: an inherent physical timescale associated with the optimal driving process.
@@
$\begin{split}-\dot{\psi}_{E}&=-\dot{\lambda}_{E}\>\partial_{\lambda_{E}}\psi_{E},\\ \langle\dot{\Sigma}_{E}^{ex}\rangle&=\sum_{\{i\}}d_{t}\rho_{i}\log{\left(\frac{\rho_{i}}{\rho_{i}^{E}}\right)}=-d_{t}D_{E}^{KL},\\ \langle\dot{\Sigma}^{hk}\rangle&=\sum_{\{\gamma^{\rightleftharpoons}\}}T_{\gamma}^{\perp}F_{\gamma}\sinh{\left(\frac{F_{\gamma}}{2}\right)}.\end{split}$ (2)
These are quasi-static work, excess, and housekeeping EPR, respectively. They physically correspond to the driving of the functional energy $E$ through a set of external control parameters $\{\lambda_{E}\}$, the relaxation toward the Boltzmann distribution dictated by $E$, and the sustained dissipative transition currents due to non-conservative forces $\{F_{\gamma}\}$, respectively. $\dot{\psi}_{E}$ and $\dot{\Sigma}_{E}^{ex}$ are the boundary terms in the control parameter space and the state space, respectively [5, 49]. $\dot{\psi}_{E}$ is integrated over time to obtain the free-energy difference between the initial and final control parameters, $\psi_{E}(\lambda_{E}^{inl})-\psi_{E}(\lambda_{E}^{fnl})$. Similarly, the KL divergence ($D_{E}^{KL}$) quantifies the statistical distance of the instantaneous state-space distribution from the equilibrium Boltzmann distribution, $\rho_{i}^{E}=e^{-E_{i}+\psi_{E}}$. Thus, a time-integrated excess EPR yields an excess EP, which is the difference between the initial and final state-space distributions and is defined with respect to the initial and final control parameters, $\Sigma_{E}^{ex}=-\Delta_{0}^{\tau}D_{E}^{KL}$. It is simplified to $\Sigma_{E}^{ex}=D_{E}^{KL}(\lambda_{E}^{inl},t=0)-D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)$, with $D_{E}^{KL}(\lambda_{E}^{inl},t=0)=\sum_{\{\rho_{i}\}}\rho_{i}^{inl}\log{\left(\rho_{i}^{inl}/\rho_{i}^{E}(\lambda_{E}^{inl})\right)}$ and $D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)=\sum_{\{\rho_{i}\}}\rho_{i}^{fnl}\log{\left(\rho_{i}^{fnl}/\rho_{i}^{E}(\lambda_{E}^{fnl})\right)}$. The mean housekeeping EPR $\langle\dot{\Sigma}^{hk}\rangle$ is a non-quadratic function of the non-conservative driving force $F_{\gamma}$ multiplied by the equilibrium traffic $T_{\gamma}^{\perp}$ (the direction orthogonal to the external driving). $T_{\gamma}^{\perp}$ quantifies the scaled equilibrium diffusion constant defined for $\gamma^{\rightleftharpoons}$ due to the equilibrium thermodynamic activity of the states $\rho_{i}$ and $\rho_{j}$. For example, for MJP/ideal CRN, $T_{\gamma}^{\perp}=\rho_{i}+\rho_{j}$, and for interacting CRN, $T_{\gamma}^{\perp}=\rho_{i}e^{\mu_{i}^{int}}+\rho_{j}e^{\mu_{j}^{int}}$, where $\mu_{i}^{int}$ is the chemical potential of $\rho_{i}$ attributed to non-ideal interactions of $\rho_{i}$ [5, 50].
@@ 3.2 The physical interpretation of kinks and the role of boundary conditions
We elaborate on the physical interpretation of kinks within our GFTOC framework. In the small driving-time limit, the fundamental principle of (ST)—the timescale separation between the environment and system degrees of freedom—is violated. This principle normally ensures that the system instantaneously relaxes to the control parameters imposed by the environment. Simultaneously, the boundary conditions enforce a constraint of reaching the final control parameters within the given finite time $\tau$ from the specified initial control parameters. Hence, for small $\tau$, due to violation of the speed-limit bound eq.˜13, the environment cannot instantaneously impose the required control parameters on the system.
Kinks solve this issue, as they reduce the geodesic distance for the finite-time driving process, mathematically expressed as $|\mathcal{G}(A_{\alpha}^{i*})-\mathcal{G}(A_{\alpha}^{f*})|<|\mathcal{G}(A_{\alpha}^{i})-\mathcal{G}(A_{\alpha}^{f})|$. This trade-off between the rush to reach the final point along the geodesic and the inability to do so within time $\tau$ produces kinks. The thermodynamic cost of each kink is $\Sigma_{bnd}^{*}/2$. The system undergoes a ‘thermodynamic shock’ at the initial and final points, generating a spontaneous discontinuous change in the control parameters, which induces a global transition of the system’s state. Within our framework, $\Sigma_{bnd}^{*}$ and $\Sigma_{bulk}^{*}$ can be interpreted analogously to heat and work, respectively: heat corresponds to the instantaneous discrete cost of a jump, and work to the continuous driving of control parameters. To our knowledge, this provides a novel thermodynamic understanding of the heat–work dual manifestation of the driving EP in finite-time optimal driven far-from equilibrium processes, which is .
@@ 3.6 Comparison and differences to model-specific known results
This contradicts its formulation of constrained and unconstrained optimization for the initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, respectively. Hence, mathematically, this is interpreted as a boundary term artefact because of the inability of assigning a thermodynamic cost to kinks within their framework. In contrast, within the minimum action formalism, a thermodynamic cost is inherently assigned for kinks, and hence unconstrained optimization is implemented for both initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, signified by a double optimization problem in eq.˜14. This implies our optimization framework uses minimum assumptions, equivalently assuming the least information about optimal protocols. This allows us to implement an unconstrained optimization problem in eq.˜14 that searches/accesses a wider control parameter space for the optimal protocol.
@@ 6 Conclusion and Outlook
In the second step, we investigate the possibility of discontinuous endpoint jumps in the optimal protocol by relaxing the assumption of slow driving. We utilize the slow-driving geodesic obtained to compute the finite-time geodesic; our analysis reveals an exact mapping between the slow-driving and finite-time optimal control problem. This strong and important result allows us to extend the framework of thermodynamic geometry developed for slowly driven systems to systems driven in any finite time. Importantly, due to the ‘minimum action principle’, a thermodynamic cost is assigned to the endpoint jumps; this instantaneous dissipation cost at the endpoints is termed a ‘thermodynamic shock’. We analytically prove that it is a model-independent generic mechanism for far-from-equilibrium (fEQ) systems, which reduces the driving dissipation cost. Within our framework, discontinuous endpoint jumps in the optimal protocol are a physical manifestation of a ‘thermodynamic shock’, implying that the optimal driven dynamics are constrained by thermodynamics.

### v2

@@ 1 Introduction
Here, based on MinAP [48, 49], we formulate the generalized finite-time optimal control (GFTOC) framework valid for any discrete-state fEQ system, by exploiting the Riemannian geometric structure in the control parameter space. We compute the optimal finite-time driving protocols that minimize the total driving EP, which exhibit the ‘kinks’. GFTOC unifies the optimal control of slowly driven and finite-time processes, as we prove an exact mapping between them. Importantly, since MinAP assigns a thermodynamic cost for any transition path realization, a thermodynamic cost for sustaining the ‘kinks’ is quantified (a boundary term in driving time), namely ‘thermodynamic shock’. Thereby, the ‘kinks’ reduce the thermodynamic cost of driving (a bulk term in driving time), since the ‘kinks’ reduce the distance to be travelled in the control parameter space. The thermodynamic interplay between the boundary and the bulk terms of driving EP gives rise to the formation of the ‘kinks’ as a dynamic consequence. The dual manifestation of the bulk-boundary term of driving EP has a physical interpretation analogous to the work-heat, but formulated here for the finite-time optimal process in ST. We show that the dynamic counterpart of this thermodynamic phenomena is dictated by the restoration of the ‘finite-time speed limit’: an inherent physical timescale associated with the optimal driving process.
@@
$\begin{split}-\dot{\psi}_{E}&=-\dot{\lambda}_{E}\>\partial_{\lambda_{E}}\psi_{E},\\ \langle\dot{\Sigma}_{E}^{ex}\rangle&=-\sum_{\{i\}}d_{t}\rho_{i}\log{\left(\frac{\rho_{i}}{\rho_{i}^{E}}\right)}=-d_{t}D_{E}^{KL},\\ \langle\dot{\Sigma}^{hk}\rangle&=\sum_{\{\gamma^{\rightleftharpoons}\}}T_{\gamma}^{\perp}F_{\gamma}\sinh{\left(\frac{F_{\gamma}}{2}\right)}.\end{split}$ (2)
These are quasi-static work, excess, and housekeeping EPR, respectively. They physically correspond to the driving of the functional energy $E$ through a set of external control parameters $\{\lambda_{E}\}$, the relaxation toward the Boltzmann distribution dictated by $E$, and the sustained dissipative transition currents due to non-conservative forces $\{F_{\gamma}\}$, respectively. $\dot{\psi}_{E}$ and $\dot{\Sigma}_{E}^{ex}$ are the boundary terms in the control parameter space and the state space, respectively [5, 49]. $\dot{\psi}_{E}$ is integrated over time to obtain the free-energy difference between the initial and final control parameters, $\psi_{E}(\lambda_{E}^{inl})-\psi_{E}(\lambda_{E}^{fnl})$. Similarly, the KL divergence ($D_{E}^{KL}$) quantifies the statistical distance of the instantaneous state-space distribution from the equilibrium Boltzmann distribution, $\rho_{i}^{E}=e^{-E_{i}+\psi_{E}}$. Thus, a time-integrated excess EPR yields an excess EP, which is the difference between the initial and final state-space distributions and is defined with respect to the initial and final control parameters, $\Sigma_{E}^{ex}=-\Delta_{0}^{\tau}D_{E}^{KL}=D_{E}^{KL}(\lambda_{E}^{inl},t=0)-D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)$, with $D_{E}^{KL}(\lambda_{E}^{inl},t=0)=\sum_{\{\rho_{i}\}}\rho_{i}^{inl}\log{\left(\rho_{i}^{inl}/\rho_{i}^{E}(\lambda_{E}^{inl})\right)}$ and $D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)=\sum_{\{\rho_{i}\}}\rho_{i}^{fnl}\log{\left(\rho_{i}^{fnl}/\rho_{i}^{E}(\lambda_{E}^{fnl})\right)}$. The mean housekeeping EPR $\langle\dot{\Sigma}^{hk}\rangle$ is a non-quadratic function of the non-conservative driving force $F_{\gamma}$ multiplied by the equilibrium traffic $T_{\gamma}^{\perp}$ (the direction orthogonal to the external driving). $T_{\gamma}^{\perp}$ quantifies the scaled equilibrium diffusion constant defined for $\gamma^{\rightleftharpoons}$ due to the equilibrium thermodynamic activity of the states $\rho_{i}$ and $\rho_{j}$. For example, for MJP/ideal CRN, $T_{\gamma}^{\perp}=\rho_{i}+\rho_{j}$, and for interacting CRN, $T_{\gamma}^{\perp}=\rho_{i}e^{\mu_{i}^{int}}+\rho_{j}e^{\mu_{j}^{int}}$, where $\mu_{i}^{int}$ is the chemical potential of $\rho_{i}$ attributed to non-ideal interactions of $\rho_{i}$ [5, 50].
@@ 3.2 The physical interpretation of kinks and the role of boundary conditions
We elaborate on the physical interpretation of kinks within our GFTOC framework. In the small driving-time limit, the fundamental principle of ST—the timescale separation between the environment and system degrees of freedom—is violated. This principle normally ensures that the system instantaneously relaxes to the control parameters imposed by the environment. Simultaneously, the boundary conditions enforce a constraint of reaching the final control parameters within the given finite time $\tau$ from the specified initial control parameters. Hence, for small $\tau$, due to violation of the speed-limit bound eq.˜13, the environment cannot instantaneously impose the required control parameters on the system.
Kinks solve this issue, as they reduce the geodesic distance for the finite-time driving process, mathematically expressed as $|\mathcal{G}(A_{\alpha}^{i*})-\mathcal{G}(A_{\alpha}^{f*})|<|\mathcal{G}(A_{\alpha}^{i})-\mathcal{G}(A_{\alpha}^{f})|$. This trade-off between the rush to reach the final point along the geodesic and the inability to do so within time $\tau$ produces kinks. The thermodynamic cost of each kink is $\Sigma_{bnd}^{*}/2$. The system undergoes a ‘thermodynamic shock’ at the initial and final points, generating a spontaneous discontinuous change in the control parameters, which induces a global transition of the system’s state. Within our framework, $\Sigma_{bnd}^{*}$ and $\Sigma_{bulk}^{*}$ can be interpreted analogously to heat and work, respectively: heat corresponds to the instantaneous discrete cost of a jump, and work to the continuous driving of control parameters. To our knowledge, this provides a novel thermodynamic understanding of the heat–work dual manifestation of the driving EP in finite-time optimal driven far-from equilibrium processes.
@@ 3.6 Comparison and differences to model-specific known results
This contradicts its formulation of constrained and unconstrained optimization for the initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, respectively. Hence, mathematically, this is interpreted as a boundary term artefact because of the inability of assigning a thermodynamic cost to kinks within their framework. In contrast, within MinAP formalism, a thermodynamic cost is inherently assigned for kinks, and hence unconstrained optimization is implemented for both initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, signified by a double optimization problem in eq.˜14. This implies our optimization framework uses minimum assumptions, equivalently assuming the least information about optimal protocols. This allows us to implement an unconstrained optimization problem in eq.˜14 that searches/accesses a wider control parameter space for the optimal protocol.
@@ 6 Conclusion and Outlook
In the second step, we investigate the possibility of discontinuous endpoint jumps in the optimal protocol by relaxing the assumption of slow driving. We utilize the slow-driving geodesic obtained to compute the finite-time geodesic; our analysis reveals an exact mapping between the slow-driving and finite-time optimal control problem. This strong and important result allows us to extend the framework of thermodynamic geometry developed for slowly driven systems to systems driven in any finite time. Importantly, due to the ‘minimum action principle’, a thermodynamic cost is assigned to the endpoint jumps; this instantaneous dissipation cost at the endpoints is termed a ‘thermodynamic shock’. We analytically prove that it is a model-independent generic mechanism for far-from-equilibrium (fEQ) systems, which reduces the driving dissipation cost. Within our framework, discontinuous endpoint jumps in the optimal protocol are a physical manifestation of a ‘thermodynamic shock’, implying that the optimal driven dynamics in finite-time are constrained by thermodynamics.